In which situation is the real interest rate highest?

A) The nominal interest rate is 25% and the inflation rate is 30%

B) The nominal interest rate is 2% and the inflation rate is 1%

C) The nominal interest rate is 8% and the inflation rate 5%

D) The nominal interest rate is 11% and the inflation rate 9%

Please provide explanation thanks
Answer #1
1 + nominal interest rate = (1+ real interest rate)*(1+ inflation)
A) Nominal interest rate 25%
Inflation 30%
1+.25 = (1+real interest rate)*(1+.30)
(1+real interest rate) = 1.25/1.30
1+real interest rate = .961538
real interest rate = -.03846.
The real interest rate is -3.85%.
B) Nominal interest rate 2%
Inflation 1%
1+.02 = (1+real interest rate)*(1+.01)
(1+real interest rate) = 1.02/1.01
1+real interest rate = 1.009901
real interest rate = .009901
The real interest rate is .99%.
C) Nominal interest rate 8%
Inflation 5%
1+.08 = (1+real interest rate)*(1+.05)
(1+real interest rate) = 1.08/1.05
1+real interest rate = 1.028571
real interest rate = .028571
The real interest rate is 2.86%.
D) Nominal interest rate 11%
Inflation 9%
1+.11 = (1+real interest rate)*(1+.09)
(1+real interest rate) = 1.11/1.09
1+real interest rate = 1.018349
real interest rate = .018349
The real interest rate is 1.83%.
The real interest rate is highest in C) that is 2.86%.
The real interest rate is highest in C
where the nominal interest rate is 8%
and the inflation is 5%.